pooja195 (pooja195):

Kara is playing a game where she flips a coin 3 times. She wins and loses different amounts of money based on the outcomes. What is the EXPECTED VALUE for the game? Question asked by : @Malawah (the user could not post questions for some reason)

OpenStudy (malawah):

Thanks pooja

OpenStudy (joshisawesome):

Pooja!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! the bewst mod ever

OpenStudy (malawah):

Here is the graph I was not able to attach @pooja195

OpenStudy (518nad):

so what is the proabability of each of those cases

OpenStudy (518nad):

@malawah are u there

OpenStudy (malawah):

Oh yes I just got your notification the theoretical probability of landing head or tails is .5

OpenStudy (518nad):

yes thats right i mean from the table you are given each of these cases is 1/8 th of the possible outides

OpenStudy (518nad):

outcomes*

OpenStudy (malawah):

i know the formula of finding expected value is (x1*p1)+(x2*p2)... and so on however i am confused on the X part since they are numbers do I multiply the x's by 0.5^3 and then continue with the formula

OpenStudy (518nad):

so u have money set = {10,3,0,-3,10,-3,0,0} each of these numbers is as likely

OpenStudy (518nad):

um i suggest you understand this formula

OpenStudy (518nad):

so the average of this set is the average value of the outcomes

OpenStudy (518nad):

{10,3,0,-3,10,-3,0,0} avg of set = (10+3+0+-3+10+-3+0+0)/8 17/8=2+1/8

OpenStudy (518nad):

i have to go now, im sure @imqwerty , can explain to you about expected value

OpenStudy (518nad):

ask him what it means and how to find it

OpenStudy (malawah):

Thanks 518 I cant thank you enough

imqwerty (imqwerty):

hello

OpenStudy (malawah):

hi :)

imqwerty (imqwerty):

so after the coins are tossed we can get different outcomes there can be 8 different outcomes as given in the list now we need to find how much money will we end up with probability of an outcome \(\times\) money received = total money we get

OpenStudy (malawah):

My computer is lagging i saw the beginning of your explanation what should I do?

OpenStudy (malawah):

I mean now i cant see it anymore im sorry

imqwerty (imqwerty):

just keep following and ask if you get stuck anywhere =]

imqwerty (imqwerty):

there are 8 possibilities the probability to get the first any of those 8 possibilities is 1/8 so the total money will be this- \(\frac{1}{8} \times 10+\frac{1}{8} \times 3+\frac{1}{8} \times 0+\frac{1}{8} \times -3+\frac{1}{8} \times 10+\frac{1}{8} \times -3+\frac{1}{8} \times 0+\frac{1}{8} \times 0\)

OpenStudy (malawah):

ooooooooooooooooooooo i understand now thanks qwerty :)

imqwerty (imqwerty):

yw :)

jabez177 (jabez177):

@Malawah, you must have the Grammarly extension installed? :) In order to post a question, that has to be uninstalled or disabled. Afterwards, you can enable it again. :)

OpenStudy (malawah):

Thanks @jabez177 UR A LIFE SAVER

OpenStudy (joshmorris):

it is 2.125

OpenStudy (callielovesyhuu):

Expected value is the sum of the probability of each even times the outcome of the event. So she wins $10 with probability 2/8, $3 with probability 1/8, $0 with probability 3/8, and loses $3 with probability 2/8. So, E = 10(2/8) + 3(1/8) + 0(3/8) - 3(2/8) =5/4 + 3/8 - 3/4 =7/4 = $1.75